Find the length of a piece of nichrome wire 0.5 mm in diameter which has a resistance of 30.52 Ω. The specific resistance of nichrome wire is 109 µΩcm?


1
Expert's answer
2021-03-02T18:07:59-0500

We can find the length of a piece of nichrome wire from the formula:


"R=\dfrac{\rho L}{A},""L=\dfrac{RA}{\rho}=\dfrac{R\dfrac{\pi d^2}{4}}{\rho},""L=\dfrac{30.52\ \Omega\cdot\dfrac{\pi\cdot(0.5\cdot10^{-3}\ m)^2}{4}}{109\cdot10^{-8}\ \Omega\cdot m}=5.5\ m."